Country A Country B World QDA = 100-10P QDB = 160-20P QDW = QSA = -40+10P QSB = -60+20P QSW = Find the equilibrium price and quantity for the world, and then fill in the table below with results after trade in widgets has taken place.
Answer #1

World demand=QdA+QdB=100-10p+160-20p=260-30p

World supply=QsA+QsB=-40+10p-60+20p=-100+30p

World equilibrium is at where world demand Equal to world supply.

260-30p=-100+30p

360=60p

P=360/60=6

World EQUILIBRIUM QUANTITY=260-30*6=80

QdA=100-10*6=40. ,QsA=-40+10*6=20

Country A will import 20 units .

QdB=160-20*6=40. QsB=-60+20*6=60

Country B will export 20 units.